Psychology 407 Assignment #10 & #11

Assignment #10 Hays — Chapter 19; 4 (sign test); 6 (Mann-Whitney); 8 (Kruskal-Wallis);

10 (Wilcoxen); 12 (Friedman)

Assignment #11 Hays — Chapter 19; 20 (Spearman rank-order correlation); 26 (Goodman-

Kruskal gamma); 30 (Kendall’s coefficient of concordance)